# Active Visual Thinking

### v1

@@ 4.1 Experimental Setup
For the base model, we use Qwen3-VL-32B-Instruct bai2025qwen3vl to train FigR. We set the maximum number of interaction rounds to 3, the maximum generation length to 32,768 tokens, and the sampling temperature to 0.7.
@@
We train our FigR on DeepMath-103K he2025deepmath, a rigorously decontaminated, large-scale dataset of 103,000 challenging mathematical problems with verifiable answers, explicitly curated to avoid overlap with standard test benchmarks. In addition, as described in Section 3.3, we annotate each instance in DeepMath-103K with a binary label indicating whether the current problem is suitable for figure-steered reasoning. The prompt template we use is shown in Appendix B.1.
@@
  • Large Language Models (LLMs). These models take textual inputs and generate textual outputs. We include Qwen3-235B-A22B (Thinking) yang2025qwen3 and Qwen3-32B (Non-Thinking, Thinking) yang2025qwen3.
  • Large Vision-Language Models (LVLMs). These models could accept both text and images as inputs and produce textual outputs. We evaluate GLM-4.5V (108B) hong2025glm45v_glm41vthinking, MiMo-VL-7B-SFT-2508 yue2025mimo_vl, and Qwen3-VL-32B-Instruct.
  • Unified Multimodal Models (UMMs). These models also take both text and images as inputs, but can generate visual outputs. We include Bagel-7B-MoT deng2025emerging and Bagel-Zebra-CoT li2025zebra.
  • Tool-Augmented Vision-Language Models (TAVLMs). These models enhance visual reasoning by invoking predefined tools or APIs to process given images. We include two RL-based models DeepEyes zheng2025deepeyes and Chain-of-Focus zhang2025chain.
In addition, we provide a text-only RL baseline, where the base model is trained with GRPO solely on the problem–answer pairs from DeepMath-103K. This setting does not involve multi-turn rollout, image generation, or any external feedback, and the reward is computed only from the final textual output—covering both answer correctness and output format. For all baseline models, we set the maximum generation length, sampling temperatures, and other hyperparameters follow the recommended configurations from their respective papers or reports. The prompt templates are shown in Appendix B.2. Due to space constraints, for UMMs and TAVLMs, we report results on a subset of datasets in the main ablation study for clarity; full results are deferred to Appendix C.
@@
MiMo-VL-7B-SFT-2508 73.80 68.07 39.00 94.40 91.57 36.40 68.40 67.38
Qwen3-VL-32B-Instruct 73.33 66.20 43.00 93.60 84.34 37.87 67.51 66.55
Text-only RL 73.33 69.22 46.00 94.40 87.95 43.38 66.18 68.64
FigR (ours) 79.58 79.32 54.00 95.00 93.98 41.92 70.33 73.45
Gains +6.25 +13.12 +11.00 +1.40 +9.64 +4.05 +2.82 +6.90
@@ 4.2 Main Results
As shown in Table 1, FigR achieves substantial gains by incorporating visual cues through end-to-end reinforcement learning, reaching an average accuracy of 73.45% across seven mathematical benchmarks, exceeding the base model (Qwen3-VL-32B-Instruct) by 6.90% and the text-only RL baseline by 4.81%. Notably, the improvements are particularly pronounced on challenging benchmarks such as AIME 2025 and BeyondAIME. FigR significantly outperforms equivalently sized text-only models (Qwen3-32B Thinking) as well as a larger LVLM GLM-4.5V. These results indicate that purely textual chain-of-thought reasoning faces clear limitations on hard problems, whereas multimodal chain-of-thought effectively unlocks the reasoning capacity of FigR.
Figure 4: Ablation analysis across training steps. We report average response length, code count, code ratio, average code lines per executable block, and code pass rate under different settings. FigR enables more frequent and structured visual reasoning while preserving stable execution behavior, compared to ablated settings.
@@ 4.3 Ablation Study
As shown in the Table 2, we conduct ablation studies on two representative datasets: AIME 2025 and BeyondAIME, to evaluate the contributions of different components in FigR.
@@
Prompt engineering introduces structured multimodal reasoning patterns without updating model parameters. While this leads to moderate performance gains on some datasets, the absence of a learning signal prevents the model from internalizing when and how such behaviors should be applied. As a result, the induced multimodal reasoning behaviors remain unstable, which inherently limits the attainable performance gains.
@@
We evaluate two baselines that directly inject visual information(+ Bagel img and + Qwen img), which generate a diagram conditioned on the problem text by using Bagel-7B-MoT and Qwen-Image wu2025qwen, respectively. Although both baselines incorporate visual inputs, neither yields consistent improvements over the base model or text-only RL. These results indicate that passively injecting images fails to provide reliable structural guidance and may instead introduce noise. This contrast highlights the importance of active, feedback-driven visual reasoning in FigR.
@@
Both unified multimodal models and tool-augmented vision-language models exhibit consistently weak performance. Unified models are vulnerable to cascading errors when tackling highly complex problems. Tool-augmented methods are restricted to predefined operations on given images and cannot flexibly construct task-specific diagrams. Consequently, neither paradigm provides reliable and informative visual feedback for complex mathematical reasoning. Due to space limitations, their full results are deferred to Appendix C
@@
We further design two ablations to examine the contributions of FigR ’s components: (1) Removing the adaptive reward mechanism (ARM). We disable the visual-invocation reward, while keeping all other components unchanged. This ablation removes the explicit control signal that regulates when visual reasoning should be invoked. (2) Removing visual feedback. Building on (1), we further remove visual feedback during reinforcement learning. The model still performs multi-turn rollouts and executes code, but receives only textual feedback. This setting isolates the contribution of visual feedback in forming a stable multimodal reasoning loop. Both ablations lead to substantial performance degradation compared to FigR, demonstrating the effectiveness of the proposed components.
@@
Evolution of active visual thinking during training.
@@
(1) When visual feedback is removed (w/o vis), the model exhibits persistently low code usage despite retaining textual execution feedback. This behavior highlights a critical limitation: textual feedback fails to introduce structurally novel or irreducible information beyond what the model can already generate in natural language. Consequently, the model learns to bypass code execution and instead relies on increasingly verbose textual reasoning, as reflected by longer responses but persistently low code ratio.
(2) Removing the adaptive reward mechanism (w/o ARM) leads to a transient increase in early code usage. However, this behavior rapidly collapses: both code ratio and code length drop to zero as the model learns that indiscriminate code invocation does not improve outcomes. This pattern reflects an overuse–abandonment cycle in the absence of effective reward guidance.
(3) In contrast, FigR maintains a consistently high code ratio throughout training, accompanied by stable and substantially longer code blocks with a high execution pass rate. Visual feedback introduces an external representational modality that cannot be easily replicated within the language space, breaking the text-only reasoning loop. Consequently, code execution becomes a necessary and reliable mechanism for constructing and validating global structure, rather than an opportunistic or exploratory behavior.
@@ 4.4 Case Study
In contrast, FigR closes the reasoning loop by incorporating visual feedback, enabling it to observe global structural patterns and adjust subsequent reasoning steps accordingly. This leads to a more robust and interpretable solution. Overall, this case study shows that the advantage of FigR does not stem merely from executing intermediate programs, but from actively integrating visual feedback into the reasoning process. Compared with purely textual reasoning or execution with text-only feedback, active visual thinking provides complementary information that substantially improves reasoning reliability.
@@ 5 Conclusion
We introduce FigR, which integrates active visual thinking into multi-turn reasoning through end-to-end reinforcement learning. By closing the loop between reasoning and visual evidence, FigR enables models to reason over global structural properties that are difficult to infer from text alone. An adaptive reward mechanism regulates when visual reasoning should be invoked, promoting selective and reliable figure construction without unnecessary generation. Extensive experiments highlight the importance of actively incorporating visual feedback during reasoning and suggest a promising direction for interpretable multimodal reasoning.
@@
MiMo-VL-7B-SFT-2508 73.80 68.07 39.00 94.40 91.57 36.40 68.40 67.38
Qwen3-VL-32B-Instruct 73.33 66.20 43.00 93.60 84.34 37.87 67.51 66.55
@@
Text-only RL 73.33 69.22 46.00 94.40 87.95 43.38 66.18 68.64
FigR (ours) 79.58 79.32 54.00 95.00 93.98 41.92 70.33 73.45
Gains +6.25 +13.12 +11.00 +1.40 +9.64 +4.05 +2.82 +6.90
Table 3: Full experimental results (%) on seven mathematical reasoning benchmarks. Green-colored font indicates the improvement of FigR over the baseline Qwen3-VL-32B-Instruct.
Figure 6: Prompt template for the figure-steered suitability classifier.
@@ B.1 Prompt Template for Suitability Classifier
We provide a dedicated prompt template for the figure-steered suitability classifier as introduced in Section 3.3 and Section 4.1, which is used to determine whether a given problem is suitable for visual-guided reasoning. As shown in Figure 6, the template instructs the model to assess the necessity and usefulness of visual sketches based solely on the problem statement.
@@ B.2 Prompt Template for Reasoning
We adopt three different categories of prompt templates corresponding to the reasoning settings. (1) As shown in Figure 7, a multi-turn prompt with drawing feedback is used for FigR, the variant without visual feedback, and the variant without ARM. Models can iteratively generate output and receive feedback from the interpreter. (2) As shown in Figure 8, a standard single-turn prompt is adopted for text-only RL and other baselines, requiring the model to directly output the final answer without intermediate interactions. (3) For DeepEyes and Chain-of-Focus, we adopt the prompt templates recommended by the original papers.

### v2

@@ 4.1 Experimental Setup
We conduct reinforcement learning using the VeRL framework sheng2024hybridflow, with Qwen3-VL-32B-Instruct bai2025qwen3vl as the base policy. We set the maximum number of interaction rounds to 3, the maximum generation length to 32,768 tokens, and the sampling temperature to 0.7.
@@
We train FigR on DeepMath-103K he2025deepmath, a rigorously decontaminated dataset of 103,000 challenging mathematical problems with verifiable answers. Following Section 3.3, we annotate each instance with a binary suitability label indicating whether executable visual construction is likely to be beneficial. An auxiliary language model (Deepseek-V3 liu2024deepseek) is used for this annotation; prompt details are provided in Appendix B.1.
@@
  • Implicit-State Reasoning Models . These models generate text-only outputs that implicitly maintain all intermediate states. We include three LLMs: Qwen3-235B-A22B (Thinking) yang2025qwen3 and Qwen3-32B (Non-Thinking, Thinking) yang2025qwen3; and three LVLMs: GLM-4.5V (108B), Qwen3-VL-8B-Instruct, and Qwen3-VL-32B-Instruct.
  • Unconstrained Visual Generation Models. Unified multimodal models (UMMs) that generate images as part of the reasoning process, including Bagel-7B-MoT deng2025emerging and Bagel-Zebra-CoT li2025zebra.
  • Tool-Augmented Vision-Language Models (TAVLMs). Models that rely on predefined visual tools or APIs operating on given images, including DeepEyes zheng2025deepeyes and Chain-of-Focus zhang2025chain.
We additionally include a text-only RL baseline trained with GRPO on DeepMath-103K, which optimizes trajectory-level correctness and format rewards without executable visual construction or interpreter feedback. For all baseline models, we set the maximum generation length, sampling temperatures, and other hyperparameters follow the recommended configurations from their respective papers or reports. The prompt templates are shown in Appendix B.2. Due to space constraints, for UMMs and TAVLMs, we report results on a subset of datasets in the main ablation study for clarity; full results are deferred to Appendix C.
@@
Qwen3-VL-8B-Instruct 61.46 46.20 30.00 94.20 81.93 40.07 67.21 60.15
Qwen3-VL-32B-Instruct 73.33 66.20 43.00 93.60 84.34 41.18 69.73 67.34
Text-only RL 73.33 69.22 46.00 94.40 87.95 43.38 70.33 69.23
FigR (ours) 79.58 79.32 54.00 95.00 93.98 44.49 72.40 74.11
Gains +6.25 +13.12 +11.00 +1.40 +9.64 +3.31 +2.67 +6.77
@@ 4.2 Main Results
As shown in Table 1, FigR achieves substantial gains by learning to integrate executable visual construction into multi-turn reasoning. Across seven mathematical benchmarks, FigR achieves an average accuracy of 74.11%, exceeding the base policy (Qwen3-VL-32B-Instruct) by 6.77% and the text-only RL baseline by 4.88%.
Performance gains are particularly pronounced on challenging benchmarks such as AIME 2025 and BeyondAIME, where reasoning requires maintaining global geometric or structural constraints. Notably, FigR outperforms both equivalently sized text-only reasoning models (e.g., Qwen3-32B Thinking) and a larger LVLM GLM-4.5V that lacks executable visual construction. These results attribute the gains to learning a policy that selectively constructs controllable, executable visual states as intermediate representations during multi-turn reasoning.
Figure 4: Ablation analysis across training steps. We track response length, code ratio, code lines, and code pass rate under different settings. FigR exhibits sustained and structured executable visual construction with stable execution behavior, whereas removing the adaptive reward mechanism causes code usage to collapse during training.
@@ 4.3 Ablation Study
As shown in Table 2, we conduct ablation studies on two representative datasets: AIME 2025 and BeyondAIME, to evaluate the contributions of different components in FigR.
@@
Prompt engineering introduces structured multimodal reasoning patterns without updating model parameters. While this leads to moderate performance gains on some datasets, the absence of a learning signal prevents the model from internalizing when and how such behaviors should be applied. As a result, the induced drawing behaviors remain unstable, which inherently limits the attainable performance gains.
@@
We evaluate two baselines that provide passively generated diagrams as static context (+Bagel img and +Qwen img). These diagrams are generated from the problem text using Bagel-7B-MoT and Qwen-Image wu2025qwen, respectively, but are not constructed or revised during reasoning. Although these baselines expose the model to visual content, they do not consistently improve performance over the base model or text-only RL baseline. These results show that passively injected visual inputs, without executable construction or feedback integration, fail to serve as reliable intermediate states. Without a closed reasoning-execution loop, such images neither enforce global constraints nor provide actionable feedback for subsequent reasoning steps.
@@
Both unified multimodal models and tool-augmented vision-language models exhibit consistently weak performance. Unified models are vulnerable to cascading errors when tackling highly complex problems. Tool-augmented methods are restricted to predefined operations on given images and cannot flexibly construct task-specific diagrams. Consequently, neither paradigm provides reliable and informative visual feedback for complex mathematical reasoning. Due to space limitations, their full results are deferred to Appendix C.
@@
We further design two ablations to examine the contributions of FigR ’s components: (1) Removing the adaptive reward mechanism (ARM). We disable the visual-invocation reward, while keeping all other components unchanged. This ablation removes the explicit control signal that regulates when visual reasoning should be invoked. (2) Removing visual feedback. We remove only visual feedback during reinforcement learning. The model still performs multi-turn rollouts and executes code, but receives only textual feedback and its corresponding calling reward. This setting isolates the contribution of visual feedback in forming a stable reasoning loop. Both ablations lead to performance degradation compared to the full FigR, demonstrating the effectiveness of the proposed components.
@@
Evolution of Active Visual Construction During Training.
@@
(1) Removing the adaptive reward mechanism (w/o ARM) leads to a brief increase in early code usage. However, this behavior rapidly collapses: both code ratio and code length drop to zero as the model learns that indiscriminate code invocation does not improve outcomes. This pattern reflects an overuse–abandonment cycle in the absence of effective reward guidance.
(2) In contrast, FigR maintains a consistently high code ratio throughout training, accompanied by stable and substantially longer code blocks with a high execution pass rate. Executable visual feedback provides an external state that enables explicit validation of global structure, which cannot be reliably enforced through implicit textual representations alone. Consequently, visual thinking becomes a reliable mechanism for validating the whole reasoning process, rather than an opportunistic or exploratory behavior.
@@ 4.4 Case Study
In contrast, FigR closes the reasoning loop by incorporating explicit visual state, enabling it to observe global structural patterns and adjust subsequent reasoning steps accordingly. This leads to a more robust and interpretable solution. Overall, this case study shows that the advantage of FigR does not stem merely from executing intermediate programs, but from actively integrating visual construction into the reasoning process. Compared with purely textual reasoning or execution with text-only feedback, active visual thinking provides complementary information that substantially improves reasoning reliability.
@@ 5 Conclusion
We introduce FigR, which integrates executable visual construction into multi-turn reasoning via end-to-end reinforcement learning. By embedding controllable and revisable visual states within the reasoning loop, FigR enables reliable reasoning over global structural constraints that are difficult to maintain through text alone. Extensive experiments demonstrate that learning when and how to externalize intermediate structure as executable visual states significantly improves reasoning stability and accuracy on complex mathematical problems.
Limitations
While our results demonstrate the effectiveness of executable visual construction for complex reasoning, several limitations remain that suggest directions for future work. The proposed method incurs additional computational cost due to multi-turn reasoning and code execution. This overhead is inherent to approaches that externalize intermediate states, motivating future efforts toward more efficient reasoning–execution loops through compact visual representations or adaptive control strategies. Executable visual construction is particularly beneficial for problems that require maintaining global structural or relational constraints. In contrast, tasks that can be easily solved through local or purely symbolic reasoning may see less benefit from visual externalization. More precise identification and characterization of such task regimes remains an open research question. Our evaluation focuses on mathematical reasoning benchmarks that emphasize structural consistency and verifiable outcomes. Extending this paradigm to other domains, including scientific reasoning, program synthesis, and long-horizon planning, represents a natural and promising direction for future work.
@@
Qwen3-VL-8B-Instruct 61.46 46.20 30.00 94.20 81.93 40.07 67.21 60.15
Qwen3-VL-32B-Instruct 73.33 66.20 43.00 93.60 84.34 41.18 69.73 67.34
@@
Text-only RL 73.33 69.22 46.00 94.40 87.95 43.38 70.33 69.23
FigR (ours) 79.58 79.32 54.00 95.00 93.98 44.49 72.40 74.11
Gains +6.25 +13.12 +11.00 +1.40 +9.64 +3.31 +2.67 +6.77
Table 3: Full experimental results (%) on seven mathematical reasoning benchmarks compared with several competitive baselines. Green-colored font indicates improvement over the baseline Qwen3-VL-32B-Instruct.
Figure 6: Prompt template for the suitability classifier.
@@ B.1 Prompt Template for Suitability Classifier
We provide a dedicated prompt template for the suitability classifier as introduced in Section 3.3 and Section 4.1, which is used to determine whether a given problem is suitable for diagrammatic reasoning. As shown in Figure 6, the template instructs the model to assess the necessity and usefulness of visual sketches based solely on the problem statement.
@@ B.2 Prompt Template for Reasoning
We adopt three different categories of prompt templates corresponding to the reasoning settings. (1) As shown in Figure 7, a multi-turn prompt is used for FigR, the variant without visual feedback, and the variant without ARM. Models can iteratively generate output and receive feedback from the interpreter. (2) As shown in Figure 8, a standard single-turn prompt is adopted for the text-only RL baseline and other models, requiring the model to directly output the final answer without intermediate interactions. (3) For DeepEyes and Chain-of-Focus, we adopt the prompt templates recommended by their original papers.